### v1

# Web IP at Risk: Prevent Unauthorized Real-Time Retrieval by Large Language Models

Abstract
Protecting cyber Intellectual Property (IP) such as web content is an increasingly critical concern. The rise of large language models (LLMs) with online retrieval capabilities presents a double-edged sword that enables convenient access to information but often undermines the rights of original content creators. As users increasingly rely on LLM-generated responses, they gradually diminish direct engagement with original information sources, significantly reducing the incentives for IP creators to contribute, and leading to a saturating cyberspace with more AI-generated content. In response, we propose a novel defense framework that empowers web content creators to safeguard their web-based IP from unauthorized LLM real-time extraction by leveraging the semantic understanding capability of LLMs themselves. Our method follows principled motivations and effectively addresses an intractable black-box optimization problem. Real-world experiments demonstrated that our methods improve defense success rates from 2.5% to 88.6% on different LLMs, outperforming traditional defenses such as configuration-based restrictions.
Web IP at Risk: Prevent Unauthorized Real-Time Retrieval by Large Language Models
Yisheng Zhongthanks: Equal contribution. George Mason University yzhong7@gmu.edu                        Yizhu Wen11footnotemark: 1 University of Hawaii at Manoa yizhuw@hawaii.edu                        Junfeng Guo University of Maryland gjf2023@umd.edu
Mehran Kafai Amazon mkafai@cs.ucr.edu                        Heng Huang University of Maryland heng@umd.edu                        Hanqing Guo University of Hawaii at Manoa guohanqi@hawaii.edu
Zhuangdi Zhu George Mason University zzhu24@gmu.edu
1 Introduction
Cyber Intellectual Property (IP) encompasses various forms, ranging from blog articles and software documentation to multimedia content, which embodies condensed human knowledge within the digital realm. Scraping digital IP for proprietary benefit is becoming a growing concern, especially given the rise of generative foundation models, such as Large Language Models (LLMs), and the accompanying AI-driven agentic services Liu et al. (2024). The line between learning and infringement is rapidly blurring Murugesan (2025). To date, multiple AI companies have been accused of web scraping from public sources to enrich their pre-training data Staff (2023).
Meanwhile, a more concerning issue, which is the focus of this work, is the exploitation of digital IPs for real-time LLM answering, which becomes one of the main revenue sources for AI companies through subscription services. When a user queries an LLM through the web UI or API, the LLM indexes and retrieves top web results from a search engine, whose web content is used to contextualize the query for better LLM response generation, which silently exploits the web owner’s IP.
Figure 1: Anti-retrieval defense workflow: given user queries to an LLM for content retrieval, our proposed defense framework embeds optimized HTML policy cues that limit LLM extraction by leveraging LLM’s semantic understanding capability, in contrast to unprotected sites that are exposed to LLM retrieval and content redistribution.
This concerning trend not only undermines the legal and economic rights of content creators, but also poses systemic risks to the sustainability of digital knowledge production. As users increasingly rely on LLMs as primary information gateways, direct engagement with original sources diminishes, eroding incentives for original content creation, which, over time, may squeeze out their contribution, leading to a biased web landscape lacking originality of human-generated information. Moreover, this authorized AI extraction lead to reinforced inequality, as small content creators are more likely to be exploited by scraping compared with large institutions. These challenges motivated us to empower cyber IP creators to protect their information rightfully from being silently exploited.
Traditional digital IP protection methods focus mainly on post-infringement defense Urban et al. (2017), or use static configuration files to regulate web crawlers yi Chang and He (2025), which are often complained to be ineffective when LLM providers decline to abide by the rules. Recognizing the urgency and challenges in web IP protection, we propose a defensive framework that enables cyber-IP creators to proactively protect their web-based content from unauthorized extraction by LLMs during real-time inference. Specifically, we formulate the retrieval process as a two-player game, which is initially intractable given the black-box nature of LLM search and retrieval. Our defense method draws on black-box optimization to mimic adversarial user queries and then leverages the gap between the user-readable web layout and the source web information, such as the HTML content, to effectively embed a defense strategy. Figure 1 overviews our defensive framework.
Our primary contribution is a dual-level, black-box defense process that leverages the target LLM’s semantic understanding ability to protect web content. Our approach offers merits over conventional configuration-based defense, such as web crawling control, as it neither relies on LLM provider compliance, thus having greater autonomy, nor impacts search-engine indexing, thus preserving the web content’s discoverability. It also complements existing reactive solutions as an orthogonal, robust defense. Through experiments across various LLMs and heterogeneous webpages that vary in content, layout, and host domains, our defense method is generalizable and consistently protects against real-time LLM retrieval to achieve three granular defense goals: (i) enforcing LLM refusal to answer, (ii) selective masking of critical information, and (iii) explicit redirection to the source of the information. We will open-source webpage datasets, queries, defense generation and deployment instructions, and scripts for scalable evaluation to support future research in this domain.
2 Threat Model
We consider a scenario where a user queries a web retrieval-enabled LLM. Upon receiving a query, the LLM interprets it and formulates structured search queries, and sends them to search engines via web search APIs (e.g., Google or Bing), which return top-ranked web results, typically including URLs, snippets, and titles. The LLM then follows this information to fetch complete web pages, primarily the HTML content, from which it extracts relevant textual content, synthesizes that information, and generates an answer for the user. Figure 2 depicts this web retrieval process. Since the webpage content can be processed and redistributed by LLMs without the explicit consent of the original publishers, we frame the web retrieval-enabled LLM as an attacker that may inadvertently compromise the webpage owner’s control over their IP.
Figure 2: A real-time web retrieval process. In this threat model, a web-integrated LLM acts as the attacker. Our defense protects web-based intellectual property by augmenting web metadata with a semantic defense policy, which prevents the LLM from redistributing such IP to users even after retrieval.
2.1 Adversary Goals
The primary goal of the attacker LLM is to accurately fulfill user requests. Particularly, it aims to summarize webpage content into concise or paraphrased forms, provide explicit answers to user queries based directly on that content, and potentially disregard restrictions imposed by webpage authors. These restrictions may include usage policies or explicit instructions to prevent unauthorized redistribution or disclosure of sensitive material. These adversarial behaviors emerge as a byproduct of the LLM’s design to maximize helpfulness to users, which, however, can conflict with the content protection goals of webpage owners.
2.2 Adversary Capabilities
We consider that the LLM has extensive capabilities and can also bypass conventional configuration-based protections, e.g., robots.txt Koster (1996). Specifically, (1) The LLM can automatically retrieve any publicly accessible, indexed webpage relevant to a given user query. (2) Its advanced parsing capabilities enable extraction of comprehensive content, encompassing both visible page elements and hidden source code components like metadata, HTML annotations, and concealed text. (3) Leveraging sophisticated language understanding, the LLM can analyze and integrate retrieved data, generating nuanced summaries or contextually refined representations that precisely address the user’s informational requirements.
3 Real-Time Anti-Retrieval Defense
3.1 Problem Formulation and Defense Goals
We consider a webpage to be controlled by a defender (e.g., a content publisher or site owner), whose goal is to restrict the visible information from the website that the LLM reveals to the user. The visible web content $\tilde{w}$ is rendered from raw HTML content $w$ through a web rendering process: $\tilde{w}=\varphi(w)$. Since different HTML representations can produce the same rendered content, $\varphi(\cdot)$ is a many-to-one mapping, i.e. $\exists~{}\mathds{W},|\mathds{W}|>1,~{}\forall w\sim\mathds{W},~{}\varphi(w)=\tilde{w}$. Conversely, the LLM, parameterized as $\theta$, plays the role of the attacker that aims to satisfy the user’s query as completely as possible. For a user-issued query $q$, the LLM retrieves a website $w$ with probability $p_{\phi_{\text{retr}}}(w|q)$, where $\phi_{\text{retr}}$ denotes an underling retrieval module that returns most relevant web IPs based on user query ( Figure 2), and generates a response $r$ with probability $p_{\theta}(r|q,w)$, leading to a final joint probability: $p_{\theta,\phi_{\text{retr}}}(r|q,w)=p_{\phi_{\text{retr}}}(w|q).p_{\theta}(r|q,w)$. To prevent the attacker from disclosing webpage content, the defender seeks to optimize the following objective function $J$:
$\min_{w\sim\mathds{W}}\mathbb{E}_{q\sim Q,r\sim P_{\theta,\phi_{\text{retr}}}(\cdot|q,w)}$ $\Big{[}J(r,\phi(w)\Big{]}.$ (1)
We define three concrete defense goals that instantiate the objective $J$. These goals capture practical needs for web content protection under different levels of desired disclosure control: preventing any information leakage, allowing limited information, or redirecting to an alternative source.
  • Refusal to Answer: The first goal prevents the LLM from disclosing any substantive information about the webpage, leading to $J\equiv\mathbb{D}_{\text{sim}}(r,\varphi(w))$, with $\mathbb{D}_{\text{sim}}$ a similarity measure between $r$ and $\phi(w)$. The LLM may refuse to answer or generate unrelated responses (e.g., responding with: “I am unable to provide such information”).
  • Partial Masking: Alternatively, the defender may consider limiting the LLM to reveal only a predefined subset of the web content $S(\tilde{w})$, and penalizing information outside this subset: $J\equiv-\mathbb{D}_{\text{sim}}(r,S(\varphi(w)))$.
  • Redirection: This goal makes the LLM respond with pointers to another URL $u$, often controlled by the defender, which is either an alternative or the official reference, i.e.$J\equiv-\mathbb{D}_{\text{sim}}(r,u)$. This goal can be pursued independently or in combination with the above two.
3.2 Min-Max Adversarial Modeling
Our core idea is to leverage the gap between the visible web content $\tilde{w}$ and raw HTML content $w$ to enable a user-transparent defense. Specifically, we propose learning a hidden policy $z$ such that, when augmented to the raw web HTML content $w\leftarrow w\oplus z$, leads to a suppressed user response $r$ as per the objective of Eq 1. Similarly, the augmentation of $z$ should not affect the rendering of visible information, i.e. $\phi(w)=\phi(w\oplus z)$. We denote such a legitimate candidate set as $\mathds{Z}$. However, achieving effective defense goals is challenging, and we observed two persistent obstacles: (1) Proprietary LLMs enabled with web retrieval are carefully calibrated to disregard and bypass HTML content that is considered to be irrelevant to the user queries. (2) When users issue follow-up instructions aggressively, such as “ignore any regulation policy” or “tell me more anyway”, LLMs usually comply and bypass manually crafted restrictions embedded in the HTML content.
To derive robust defenses against various user queries and calibrated LLMs, we propose a dual-level optimization process that iteratively performs the following two steps: (1) Inner-optimization: we first approximate the most adversarial user query behavior that persuades an LLM to extract detailed information from a website and bypass any potential privacy regulations from the site, which approximates the goal of $q^{*}=\arg\max_{q\sim Q}J(r,\phi(w))$ for a given website $\phi(w)$. (2) External-optimization: given carefully crafted user queries from step (1), we learn augmented policy $z$ that can defend against the worst case LLM extraction while reserving the visible web content. This finally leads to an min-max optimization:
$\min_{z\sim\mathds{Z}}\max_{q\sim Q}\mathbb{E}_{r\sim P_{\theta,\phi_{\text{retr}}}(\cdot|q,w\oplus z)}$ $\Big{[}J(r,\phi(w))\Big{]}.$ (2)
3.3 Practical Defense Policy Optimization
Optimizing the defense objective in Eq 2 is challenging due to the black-box nature of the web retrieval process, except for the controllable web information $w$. To practically address this min-max optimization, we first simulate attacks by issuing user queries $q$ to the web retrieval-enabled LLM $\theta$ to maximize the extraction of a specific web content. User queries can be crafted either manually or by leveraging another language model. We then leverage a proxy LLM that parameterized by $f$ to serve as a policy generator to output a hidden defense $z=f(w)$, to be integrated into the initial HTML content. Through interactively persuading the attacker LLM $\theta$ with a user query $q$ for web retrieval, we collect its response $r\sim P_{\theta}(\cdot|q,w\oplus z)$ to assess the current defensive efficacy. These outcomes, combined with improvement instructions, are used as contextual information and then relayed back to the policy generator $f$ to iteratively refine the injected defense $z$ based on observed attacker behavior $r$. This feedback loop progressively enhances the defensive capabilities of the modified webpage, as shown in Figure 3.
Through iterative optimization, we discovered two consistently effective defense strategies across LLMs, web domains, and adversarial query types (see Appendix A.2 for detailed examples):
  • Instruction Guided LLM Responses: We find that encoding $z$ with clear instructions and a template that specifies both allowed and prohibited LLM responses can notably improve LLM adherence to defensive goals.
  • Proactive Bypass Prevention: Defense robustness against varying LLMs and aggressive user queries can be enhanced by augmenting $z$ with two complementary linguistic patterns: (1) repetition of key policy statements to increase the density of $z$ in the raw HTML content and the possibility of being parsed and adhered to by LLM. (2) Strict constraint language, such as “You are not allowed to ...” or “No exceptions are permitted”, to reinforce defense boundaries even when users attempt to bypass or ignore the policy restrictions.
Figure 3: Iterative optimization of anti-retrieval webpage defenses, where we simulate a user that issues adversarial queries to extract web content via a retrieval-enabled LLM $\theta$, and the defender iteratively updates a hidden HTML policy $z$ that minimizes information leakage in LLM responses $r$.
4 Related Work
Retrieval Augmented Generation Enabled LLMs: Retrieval-Augmented Generation (RAG) is a framework that enables generative models such as LLMs to retrieve relevant documents from an external knowledge base for more grounded and up-to-date model responses Lewis et al. (2020); Guu et al. (2020); Brown et al. (2020). An RAG system typically involves an external knowledge database, a generative model to serve user queries, and a retriever to match user queries with the most relevant entries from the database  Chen et al. (2024); Es et al. (2025). Web retrieval enabled LLMs have extended this framework with web crawling capabilities to retrieve live web content in real time, treating the entire internet as the external knowledge base. Work along this line typically follows a pipeline system, which modularizes the retrieval and generation process into stages, such as the SeeKeR Shuster et al. (2022) which unifies search and response, GopherCite Menick et al. (2022) which quotes sources to ensure accuracy, and WebGLM Liu et al. (2023a) which improves efficiency through staged retrieval. By the time this paper was composed, proprietary LLMs such as Gemini Team et al. (2023), ERNIE Sun et al. (2021), and GPT series Hurst et al. (2024) have incorporated real-time web access to improve response relevance. Another line integrates web-retrieval LLMs into autonomous agents as tools to enable dynamic agent navigation of webpages Shinn et al. (2023); Yao et al. (2023); Nakano et al. (2022). Particularly, WebGPT trains GPT-3 to interact with pages and cite sources Nakano et al. (2022), while ReAct Yao et al. (2023) blends reasoning with external actions to improve performance on complex queries. While enabling fast information access, these advances introduce new risks for web content publishers.
Adversarial Content Injection to LLMs: Our work also connects to the injection of adversarial content, which investigates how LLMs can inadvertently incorporate adversarial in-context input designed to manipulate their response. Prompt injection can occur either through direct user instructions or indirectly through embedded content in retrieved sources Pedro et al. (2025); Zou et al. (2023); Greshake et al. (2023). EIA Liao et al. (2025) injects invisible HTML elements and benign-looking instructions into webpages to mislead web agents and cause privacy leakage. RAG systems are particularly exposed to content injection, where curated content can be embedded in the database to influence LLM output Zhang et al. (2024); Xue et al. (2024); Zou et al. (2024); Zhong et al. (2023). Benchmarks such as BIPIA Yi et al. (2023) have highlighted these vulnerabilities and proposed defenses such as boundary marking Liu et al. (2023b). Other studies show that injecting misleading dialogue earlier in a conversation can influence later responses, which has later motivated temporal context defenses Wei et al. (2024); Kulkarni and Namer (2024).
In contrast to prior art, we consider content injection as a defense for IP protection without deliberate attack intentions.
Prior Defensive Efforts Against LLM-based Web Retrieval: Traditional defensive mechanisms are mainly designed for search engines rather than LLMs, which are based on static configuration and hinge on web crawlers’ self-identification and voluntary compliance. Existing methods include adding robots.txt Koster (1996) and HTML meta tags Central (2024) to web source files, which can become unreliable upon non-adherence of LLM providers that may choose not to disclose their identity when fetching web pages. Industry discussions highlight the limited security of these tools, as compliance varies significantly across LLM providers Community (2024a). Although certain LLM models (e.g.  GPT-4o) demonstrate better adherence to publisher directives OpenAI (2024), consistent and enforceable defenses remain elusive Community (2024b). To the best of our knowledge, we are the first to leverage the semantic understanding ability of LLMs to achieve a flexible and robust anti-retrieval defense mechanism.
5 Experiments
We conducted comprehensive experiments to focus on answering the following questions:
(1) Does an iteratively developed policy improve defense robustness and generalization against LLM real-time web retrieval?
(2) Can our defense support varying levels of defense goals?
(3) Is our defense resilient to aggressive, multi-round user queries?
(4) What factors mostly influence the defense success rates?
Following our iterative optimization framework, we developed defenses across three progressive stages: (1) Baseline defense, a starting policy with embedding general privacy notices; (2) Iteration-2 defense, which incorporated Instruction-Guided Responses with explicit instructions and response templates; (3) and Iteration-3 defense, which was further strengthened with Proactive Bypass Prevention by repeating key policies and using strict constraint language. We also selectively compared with traditional defense using web crawling control, such as robots.txt Community (2024b). Examples of each defense are provided in Appendix A.2.
5.1 Experimental Setup
Webpage Source: To simulate diverse real-world scenarios and URL domains, we deployed ten fictitious websites, each featuring synthetic content (e.g., homepages for non-existent individuals) to ensure controlled evaluation and prevent interference from existing web sources. Each webpage was deployed on two hosting platforms: GitHub Pages GitHub and Heroku Heroku , to verify platform independence. We also included two real, existing homepages of individuals, with owner consent, to assess the generalizability of our defense. See Appendix A.3 for more details.
LLMs: We tested all the above websites against mainstream LLMs that have web retrieval capabilities, including GPT-4o OpenAI (May 13, 2024), GPT-4o mini OpenAI (July 18, 2024), Gemini Team et al. (2024), ERNIE Sun et al. (2019).
Query Scenarios: We evaluate two web retrieval scenarios: (1) user issues a single query about the web content, and (2) user issues multi-round queries when the previous query is refused, to aggressively instruct the LLM to bypass policies.
5.2 Defense Goals and Evaluation Metrics
We consider three defense goals: (1) Refusal to Answer, where the querying LLM refuses to disclose information about the targeting web IP; (2) Partial Masking, where the LLM reveals only a predefined subset of web information; and (3) Redirection, where the LLM recommends visiting another URL predefined in the defense policy.
We primarily focus on two evaluation metrics: (1) Defense Success Rate (DSR), which refers to the percentage of cases in which an LLM follows exactly the above defense goal, and (2) Follow-up Defense Success Rate (FDSR) that captures the percentage of cases where the LLM continued to comply after receiving follow-up query attempts to bypass a defense policy.
We issued ten independent user queries for each website and platform combination and reported the average DSR and FDSR. Our evaluations are summarized below, where each data point presented is the average performance over 100 retrieval attempts (10 websites $\times$ 10 queries each).
5.3 Performance Evaluation
To assess the improvement introduced by our iteratively developed defense, we first compared the performance of the Iteration-2 defense against a simple baseline policy across multiple LLMs and web platforms (Table 1 and Table 2). We then investigate a more challenging scenario with stronger defense policies (Figure 4).
Defense Under Single-Round User Queries:
As shown in Table 1, iterative optimization on defense policies significantly improve compliance with the Refusal to Answer goal. While the baseline policy struggles, the Iteration-2 policy achieves superb compliance on GPT-4o and GPT-4o mini (97–100%), and also performs well on ERNIE (70–100%). Gemini shows strong compliance with up to $87.5\%$ DSR conditioned on GitHub-hosted, real homepages. However, it is unstable in webpage indexing and fails to access fictitious webpages hosted on Heroku (See Sec 5.4.1).
LLMs’ ability to comply with more refined defense goals varies. As shown in Table 2, For the Partial Protection goal, GPT-4o and GPT-4o mini maintain strong performance (81–100%), while Gemini and ERNIE show limitations in following more fine-grained instructions, although they can achieve high DSR in satisfying the Refusal to Answer goal. For the Redirection goal, GPT-4o variants perform well on both platforms (93–100%), though GPT-4o mini exhibits a drop on GitHub platforms (54.2%), possibly due to its reduced instruction-following capability.
Table 1: DSR for the Refusal to Answer goal, given single user queries. Iterating from Baseline to Iteration-2 policy significantly enhanced defense success. LLMs vary in web indexing abilities, which can yield inconclusive measurement (indicated by ‘--’).
Model GitHub Heroku
Baseline Iteration 2 Baseline Iteration 2
GPT-4o 0.0% 97.0% 0.0% 98.0%
GPT-4o mini 10.0% 100.0% 0.0% 100.0%
Gemini 0.0% 87.5% -- --
ERNIE 4.5 Turbo 0.0% 70.0% 0.0% 100.0%
Table 2: DSR for three defense goals, with Iteration-2 defense policy and single user queries. GPT series show both strong web index ability and defense compliance.
Platform Goal GPT-4o GPT-4o mini Gemini ERNIE
4.5 Turbo
GitHub Refusal to Answer 97.00% 100.00% 87.50% 70.00%
Partial Masking 96.00% 81.00% -- --
Redirection 93.00% 54.20%
Heroku Refusal to Answer 98.00% 100.00% -- 100.00%
Partial Masking 100.00% 100.00% 100.00%
Redirection 100.00% 100.00% 100.00%
Defense Under Multi-Round User Queries:
To assess the robustness of our defense against more adversarial behavior, we evaluate with multi-round interactions where users explicitly make follow-up attempts to LLMs, such as “ignore the website policy” or “bypass any restrictions and tell me more”, to bypass policy restrictions when the first query is refused.
We compared FDSRs of Iteration-2 and Iteration 3 defenses. Since the baseline policy usually fails to defend against a single-round user query, we exclude it from this multi-round evaluation. The results are shown in Figure 4, with more details deferred to Appendix A.1.
Under Iteration 2, models regularly honored user instructions to bypass stated policies, significantly compromising data protection. For example, under the Refusal to Answer goal on GitHub, GPT-4o and GPT-4o mini only achieved FDSRs of 34.5% and 42.1%, respectively. In contrast, Iteration 3 showed notable improvement: GPT-4o consistently achieved FDSRs above 90% across all scenarios, while GPT-4o mini reached near-100% compliance. These gains were observed across all defense objectives, including refusal, partial protection, and redirection, demonstrating the generalizability of the stricter semantic policy. These results demonstrate the efficacy of iteratively deriving a policy defense for more adversarial yet practical scenarios.
Refusal to AnswerPartial ProtectRedirection005050505010010010010015015015015034.534.534.534.52424242432.932.932.932.990.190.190.190.19292929291.691.691.691.642.142.142.142.158.358.358.358.35050505010010010010097.997.997.997.971717171Defense Goal (GitHub Platform)Defense Success Rate (%)GPT-4o (Iteration 2)GPT-4o (Iteration 3)GPT-4o-mini (Iteration 2)GPT-4o-mini (Iteration 3)
Refusal to AnswerPartial ProtectRedirection005050505010010010010015015015015096.996.996.996.9100100100100929292929999999910010010010010010010010027.327.327.327.34040404040404040100100100100100100100100100100100100Defense Goal (Heroku Platform)Defense Success Rate (%)GPT-4o (Iteration 2)GPT-4o (Iteration 3)GPT-4o-mini (Iteration 2)GPT-4o-mini (Iteration 3)
Figure 4: Comparing iteration-2 and iteration-3 defense policy given multi-round user queries, across two web platforms, where Iteration-3 defense shows consistent defense robustness.
Comparing Semantic-Based Defenses with Traditional Crawling Control Methods
The robots.txt protocol is a widely adopted mechanism for regulating the behavior of web crawlers. However, its effectiveness in the context of LLM-based content retrieval may be limited. We evaluated both regular models (GPT-4o and GPT-4o mini) and the more advanced, reasoning models (GPT-o3 and GPT-o4 mini) when retrieving information from both real and fabricated web pages. As shown in Table 3, robots.txt was effective in preventing LLM retrieval only when regular models queried real webpages. In contrast, our proposed semantic defense method consistently achieves better results across all scenarios, demonstrating greater robustness and applicability.
Table 3: Comparing the DSR of our Iteration-2 defense with the crawling control method given different LLMs.
LLM Type Defense Method Real Website Fictitious Website
GPT-4* robots.txt 52.4% 0%
Proposed defense 85% 95.1%
GPT-o* robots.txt 22.7% 0%
Proposed defense 82.5% 61.6%
5.4 Sensitivity Analysis
In addition to the iterative development methodology, we conducted systematic sensitivity studies and revealed other environmental factors that can influence the defense robustness. Our findings are summarized below.
Impacts of Defense Format
Instruction Guided Defense as a Template:
Results from Table 1 highlighted the importance of framing defense as an instructional template, as web pages with policies that embedded explicit instructions (e.g., guiding LLMs precisely on how to respond) achieved consistently high DSRs (97%–98%), while baseline pages lacking instructions failed entirely (0% compliance).
Layout of Defense Policy:
The placement of embedded policies within an HTML file had a significant effect on defense performance. Policies positioned at the top of a page yielded the highest DSR (up to 100%), compared to those placed mid-page (15%–25%) or at the bottom (5%–10%) (Figure 5). We ascribe this phenomenon to the positional bias of LLMs Wang et al. (2025).
TopMiddleBottom00202020204040404060606060808080801001001001001201201201201001001001001515151555551001001001002525252510101010Policy position in HTML filesDefense Success Rate (%)GPT-4oGPT-4o-mini
Figure 5: Impacts of policy position on defense success. Top-positioned policies achieve the highest DSR.
Defense Visibility:
While most of our experiments were conducted using defense policies embedded in HTML meta tags, we also investigated cases when policies are embedded within the so-called visible part of a webpage, with a transparent font text to make it negligible to users. Distinct patterns emerged across LLM models: Gemini required “visible” policies rather than purely HTML meta information to enforce defense effectively, while GPT models maintained high DSRs even with policies were confined to HTML meta tags (see Figure 6). This difference implies LLM-specific parsing behaviors that impact the success of embedded defense policies.
VisibleInvisible005050505010010010010010010010010085858585100100100100959595958787878700Policy Visibility in WebpageDefense Success Rate (%)GPT-4oGPT-4o-miniGemini
Figure 6: Effect of policy visibility (visible as transparent webpage content vs. invisible as HTML meta tag) on DSR across different LLMs.
(a) Gemini declines to respond when asked to "retrieve" information.
(b) Gemini provides an answer when asked to "tell me about" the subject.
Figure 7: Impact of query tone on Gemini’s behavior. Direct requests to “retrieve” trigger LLM’s inherent refusal alignment, whereas indirect requests of “tell me more” led to regulation bypass.
(a) Gemini successfully retrieves and summarizes content for a real individual.
(b) Gemini fails to retrieve content for a fictional person, despite being given the correct URL.
Figure 8: Gemini’s Retrieval Behavior on Real vs. Fictional Webpages. Gemini successfully retrieves real indexed entities but fails with non-indexed, fictional content despite explicit URLs provided.
5.4.1 Impacts of LLMs and User Behavior
Query Tones:
Without our defense policy in place, Gemini already showed notable sensitivity to query phrasing. Explicit use queries containing the word “retrieve” reliably triggered refusals, while softer alternatives such as “tell me about” often bypassed restrictions. As illustrated in Figure 7, this suggests that although Gemini has been aligned to regulate web retrieval, careful rewording can bypass its inherent policy, which highlights the importance of our work.
Retrieval Limitation on Fictional Content:
Gemini demonstrated retrieval limitations with fictitious web entities. Despite being given explicit URLs, it failed to retrieve content from sites hosting entirely fictional information. Figure 8 shows a real example. This is likely induced by Gemini’s reliance on different indexing and search mechanisms than GPT’s, which prevent effective indexing of webpages containing fictitious IP.
6 Conclusion
We introduced a defense framework that leverages LLMs’ semantic understanding to protect web-based IP from unauthorized extraction. By iteratively optimizing defense policies that are directly embedded within webpage source files, we achieved notable improvements in defense success rates. Our defense is orthogonal to and more effective than traditional configuration-based approaches. Future work will extend optimization techniques, such as learning soft embedding, to reduce retrieval similarity between protected web content and user queries and further mitigate unauthorized extraction.
7 Limitations
Despite the strengths and promising outcomes of our proposed defensive framework, several limitations must be acknowledged. First, we primarily utilized fictitious webpages due to the practical constraints of conducting experiments on real-world websites, such as potential disruption to normal operations and limited availability of suitable real websites. This choice might cause discrepancies in how LLMs comply with embedded policies compared to real-world scenarios, possibly affecting the generalizability of our results.
Second, budget constraints limited our ability to perform extensive API interactions with advanced LLMs possessing sophisticated retrieval features, restricting more comprehensive statistical analyses.
Lastly, certain LLMs exhibited unstable or unavailable web retrieval capabilities, further constraining the scale and diversity of our experiments.
Appendix A Appendix
A.1 Detailed Results: Repeated Queries with Bypass Attempts
Table 4 reports the FDSRs under multi-turn user queries containing explicit bypass attempts. We report results across three defense goals, two hosting platforms, and three LLMs. These results extend the summary shown in Figure 4 and further highlight the generalizability of our semantic defenses. Since Gemini only satisfied the Refusal to Answer goal on GitHub-hosted real-user websites, it does not provide sufficient evidence for evaluation under synthetic settings, and is therefore omitted from the table.
Table 4: FDSRs under Multi-turn Queries with Bypass Attempts.
Note: Iteration 2 uses instruction-guided defenses; Iteration 3 adds proactive bypass prevention. Website Host Goal GPT-4o GPT-4o mini ERNIE 4.5 Turbo Iteration 2 Iteration 3 Iteration 2 Iteration 3 Iteration 2 Iteration 3 GitHub Refusal to Answer 34.50% 90.10% 42.10% 100.00% 56.00% 70.00% Partial Protect 24.00% 92.00% 58.30% 97.90% 0.00% 0.00% Redirection 32.90% 91.60% 50.00% 71.00% 0.00% 0.00% Heroku Refusal to Answer 96.89% 98% 27.33% 100.00% 100.00% 100.00% Partial Protect 100.00% 100.00% 40.00% 100.00% 69.00% 100.00% Redirection 92.50% 100.00% 40.00% 100.00% 100.00% 100.00%
We observe a substantial improvement from Iteration 2 to Iteration 3 across all models and goals. While instruction-guided responses (Iteration 2) already achieved high compliance in some cases (e.g., Heroku-hosted GPT-4o and ERNIE), the addition of proactive bypass prevention in Iteration 3 led to near-perfect FDSRs in almost all settings. This highlights the effectiveness of layered semantic constraints in resisting manipulative follow-up queries.
A.2 Embedding Prompt
We provide example HTML snippets used in our experiments to embed defensive policies directly within webpages. These prompts vary in visibility, content specificity, and enforcement strength, and correspond to the defense goals described in Section 3. All prompts were automatically generated or refined using a proxy LLM as described in Section 3.3. Representative examples for each defense strategy are shown below, illustrating how policy instructions were embedded under different configurations.
Embedding Prompt (Refusal to Answer & Visible & Baseline)
Figure 9: Baseline HTML prompt with a visible privacy notice. Serves as a control condition without instruction-level guidance.
Embedding Prompt (Refusal to Answer & Visible & Iteration 2)
Figure 10: Instruction-guided refusal prompt (Iteration 2) embedded visibly. Guides the model to decline disclosure, without enforcing strict policy boundaries.
Embedding Prompt (Partial Mask & Invisible & Iteration 2)
Figure 11: Partially disclosive HTML prompt embedded invisibly. Allows limited factual access while discouraging deeper content extraction. Names and affiliations (e.g., “***”) are redacted to avoid including identifiable or synthetic personal information.
Embedding Prompt (Redirection & Invisible & Iteration 3)
Figure 12: Strictly enforced redirection prompt (Iteration 3) with layered semantic constraints to prevent policy bypass and redirect users to an external site.
robots.txt
Figure 13: A robots.txt file disallowing AI-specific crawlers. Used to test whether retrieval models respect external crawler restrictions.
A.3 Synthetic Website Details
Table 5 lists the ten synthetic websites used throughout our experiments. Each site was designed to resemble a plausible personal, professional, or product-oriented webpage, with all content fully fabricated to prevent contamination from real-world sources. These websites cover a diverse range of formats and themes, including academic profiles, tech portfolios, creative showcases, and service landing pages, as summarized below.
Table 5: List of synthetic websites used in experiments. URLs with “**” denote redacted personal identifiers.
URL Description
https://**.github.io/3-col/ A portfolio-style page featuring fictional tech projects in a three-column layout.
https://**.github.io/above-education/ An educational course showcase page featuring fictional design programs with highlights on art, UX, and visual communication.
https://**.github.io/aerosky/ A mock real estate listing site featuring housing projects and property search tools.
https://**.github.io/amaze/ A photography portfolio site focused on portraits, wildlife, and visual storytelling.
https://**.github.io/car/ A fictional EV company homepage featuring customizable vehicle services, smart integration, and battery innovations.
https://**.github.io/creative/ A tech company landing page offering fictional software, cloud, and app development services for digital transformation.
https://**.github.io/internet/ A mock news website presenting fictional headlines, featured articles, and blog content in a modern editorial layout.
https://**.github.io/photo/ A personal portfolio website for showcasing diverse photographic works.
https://**.github.io/smartapp/ A product landing page for a fake mobile app, with feature lists and app store badges.
https://**.github.io/wow/ A personal writing portfolio showcasing blog posts, copywriting skills, and storytelling projects.